molly is 4 years younger than heidi. in 3 years, heidi will be twice as old as molly. how old are they now?
Answer by Paul(988)

Heidi's age = x
Molly's age = x-4
In three years Molly's age = x-4+3
In three years Heidi's age = x+3

2x-2=x+3
x=5
5-4=1

Hence, Heidi is 5 years old and Molly is 1 year old.
Paul.